OpenStudy (anonymous):

(22x^5y^3z/16x^3z^5) times (8y^2z^3/11x^3)

OpenStudy (anonymous):

I would simplify what can be simplified and then multiply them together. When multiplying same bases add exponents.

OpenStudy (anonymous):

\[22x^5y^3z/ 16x^3z^5 = (22/16)x^2y^3z^{-4}\] and \[8y^2z^3/11x^3 = (8/11)y^2z^3x^-3\]that multiplied is: \[(x^{-1}y^5z^{-1} = y^5/(xz)\] Just add the exponents for each variable, then do the multiplication of the coefficients(which ends up being 1)

OpenStudy (anonymous):

thankss so much!

OpenStudy (anonymous):

\[7/(3-\sqrt{2})\]

OpenStudy (anonymous):

What about it?

OpenStudy (anonymous):

.................how do you do it.?

OpenStudy (anonymous):

Rationalize it or what?

OpenStudy (anonymous):

in order to get rid of teh radical in the denominator you have to multiply the top and bottom by the conjugate